one side of the triangle is 2 meters shorter than the base, the other side is 3 meters longer than twice the base. IF the perimeter is 37 meters, find the length of the base
Answer by ReadingBoosters(3246)

x = one side
y = other side
z = base
x+y+z=37
x = z-2
y = 3+2z
Substitute x and y in the equation
(z-2)+(3+2z)+z=37
4z +1 = 37
4z = 36
z = 9 m is the base